OpenStudy (anonymous):

For her phone service, Leila pays a monthly fee of $27, and she pays an additional .05 per minute of use. The least she has been charged in a month is $91.30 What are the possible numbers of minutes she has used her phone in a month? Use m for minutes and solve for m.

OpenStudy (anonymous):

27+0.05m=91.30 solve for m

OpenStudy (anonymous):

0.05m=91.30-27 0.05m=64.30 m=64.30/0.05

OpenStudy (anonymous):

Wouldn't I have to use an inequality for this?

OpenStudy (anonymous):

you can, so that would be the least amount

OpenStudy (anonymous):

Ok thank you!